3. The competitive market demand and supply for TV in USA are; Qd=100–20P and Qs=20+20P.
a.Sketch the demand and supply function for the STV industry.
b.Find the consumer surplus and producer surplus.
c.If the free market price is too high for the ordinary American to pay. What would happen if they set a price ceiling at $1.5.
d.If the free market price is too low to enable producers to earn a fair rate of return. What would happen if the government agrees with him and buys S TV from suppliers at a price floor of $4?
e.Find the cost of the price floor in (e) above?
f.If income of consumers rises by $10. Find the new equilibrium price and quantity?
g.If a fall in wages affects the market supply curve from Q=2+2P to Q=6+2P, how would that affect the market-clearing price and quantity?
h.find the price elasticity of demand and show whether the firm should increase price or reduce price to maximize revenue? Interpret your answer in (i) above.
1
Expert's answer
2020-11-12T17:22:58-0500

a. The demand and supply curves for the STV industry intersect in equilibrium point, where Qd = Qs, so:

"100 \u2013 20P = 20 + 20P,"

P = 2, Q = 60 units.

b. The consumer surplus is:

"CS = 0.5\u00d7(5 - 2)\u00d760 = 90."

The producer surplus is:

"PS = 0.5\u00d7(60 + 20)\u00d72 = 80."

c. If they set a price ceiling at $1.5, then the shortage of TV will occur.

d. If the government buys S TV from suppliers at a price floor of $4, then there will be no surplus.

e. The cost of the price floor in (e) above is 4S.

f. If income of consumers rises by $10, then both equilibrium price and quantity will increase.

g. If a fall in wages affects the market supply curve from Q=2+2P to Q=6+2P, then both market-clearing price and quantity will decrease.

h. The price elasticity of demand is:

"Ed = -20\u00d72\/60 = -2\/3," so the demand is inelastic.

The firm should increase price to maximize revenue.